### v1

@@ 3.3 Triple Generating Module
In this module, the 2D sentence representation $\mathbf{T}^{so}_{i,j}$ is fed into a two-layer perceptron to learn a table feature for each relation candidate. The table feature for a relation $r$ is denoted as $\mathbf{T}^{r}$. Each item in $\mathbf{T}^{r}$ represents the label feature for a token pair. Specifically, for a token pair $(x_{i},x_{j})$, we represent its label feature as $\mathbf{T}^{r}(i,j)$, which is computed by Equation (10).
@@
where $W\in\mathbb{R}^{D_{h}\times d}$, $b$ are trainable weight and bias. $W_{r}\in\mathbb{R}^{d\times 4K}$, $b_{r}$ are trainable weight and bias. Here, $K$ signifies the number of relations, 4 signifies the number of classification tags. The function $drop(\cdot)$ denotes the dropout strategy (Srivastava et al., 2014) utilized for mitigating overfitting.
Next, by utilizing the table features $\emph{TF}_{r}$, all triples can be generated. Specifically, for every relation, the method outlined in Eq. 11 is employed to populate its corresponding table.
$\hat{\operatorname{table}_{r}}(i,j)=\operatorname{softmax}\left(\emph{TF}_{r}(i,j)\right)$ (11)
@@
Finally, the filled tables are decoded. All triples are deduced using the following method. The subject can be obtained through splicing from ‘B-E’ to ‘E-E’. The object can be obtained through splicing from ‘B-B’ to ‘B-E’. They result in the acquisition of the relationship triplet. It is noteworthy that the two paired entities share a common ‘B-E’.
@@
$\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j=1}^{N}\log\left(\hat{\operatorname{table}_{r}}(i,j)=\hat{\operatorname{gold}_{r}}(i,j)\right)$
where $\hat{\operatorname{gold}_{r}}(i,j)$ denotes the gold tag obtained from annotations.
@@ 4.1 Experimental Settings
Train Valid Test Relation Normal SEO EPO SOO I=1 I=2 I=3 I=4 I>5 Triples
@@
In these datasets, both NYT and WebNLG have two distinct versions. Specifically, $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ are annotated with the tail tokens of a given entity. $\mathrm{NYT}$ and $\mathrm{WebNLG}$ are annotated with the whole span of an entity. Table 1 presents the statistical information of these datasets.
@@
To evaluate the performance of our models, we utilized Partial Match and Exact Match for the NYT, NYT10, NYT11 and WebNLG datasets. Partial Match considers a triple (subject, relation, object) to be correct only if the relation, as well as the tail tokens of both subject and object, is accurate. On the other hand, Exact Match deems a triple correct only if the entire span of both entities and the relation precisely match.
@@
Implementation Details: in order to ensure fair comparison, we utilize the Bert-base-cased model  [^1] Available at https://huggingface.co/bert-base-cased, consisting of 12 Transformer blocks with a hidden size of $D_{h}=768$. For the WebNLG dataset, a batch size of 6 is set, whereas for other datasets, the batch size is configured to be 5. We configure the local consolidation module as ‘$[C-A_{r}-R-V]$’ and optimize all parameters using the Adam (Kingma and Ba, 2014) with the learning rate of $1\times 10^{-5}$. In equation 10, a hidden layer $d=3\times D_{h}$ d is utilized for the NYT and $\mathrm{NYT}^{*}$ dataset, whereas for the remaining datasets, a hidden layer $d=2\times D_{h}$ is employed. Additionally, the maximum sentence length for input is set at 100.
@@ 4.2 Main Results
In this experiment, our model is compared with fourteen related works. In these models, GraphRel, NovalTagging, CopyRE, ETL-span, and MultiHead utilized LSTM networks as the sentence encoder. Other models applied pre-trained BERT models for generating token representations. The performance is evaluated in terms of the Partial Match and the Exact Match. The result is presented in Table 2.
@@
Model $\mathrm{PartialMatch}$ $\mathrm{ExactMatch}$
@@
PRGC (Zheng et al., 2021) $\mathbf{93.3}$ 91.9 92.6 94.0 92.1 93.0 $\mathbf{93.5}$ 91.9 92.7 89.9 87.2 88.5
@@
BiRTE (Ren et al., 2022) 92.2 $\mathbf{93.8}$ $\underline{93.0}$ 93.2 94.0 93.6 91.9 $\mathbf{93.7}$ 92.8 89.0 89.5 89.3
@@
BTDM (Zhang et al., 2023b) $\underline{93.0}$ 92.5 92.7 94.1 93.5 93.8 93.1 92.4 92.7 90.9 90.1 90.5
BitCoin (He et al., 2023) 92.9 92.8 93.1 $\mathbf{94.4}$ $\underline{94.5}$ $\underline{94.4}$ 93.1 92.6 $\underline{92.8}$ $\underline{91.9}$ $\underline{90.5}$ $\underline{91.2}$
Ours 92.9 $\underline{93.3}$ $\mathbf{93.1}$ $\underline{94.1}$ $\mathbf{96.0}$ $\mathbf{95.0}$ 93.2 $\underline{92.6}$ $\mathbf{92.9}$ $\mathbf{92.2}$ $\mathbf{91.3}$ $\mathbf{91.7}$
Model $\mathrm{PartialMatch}$ $\mathrm{ExactMatch}$
@@
As Table 2 shown, compared with other models, our model exhibits competitive performance across all datasets. In the $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ datasets, relation triples solely be annotated with the final word of a given entity. The improvement of our model suggests that the bi-consolidating model is successful in learning global features relevant to a relation instance. On the other hand, the improvement in the $\mathrm{NYT}$ and $\mathrm{WebNLG}$ also indicates that the bi-consolidating model has the ability to correctly extract the whole span of entities, which heavily depends local features of a sentence. As shown in Table 1, these datasets usually contain relation triples with the serious semantic overlapping problem. The outcome demonstrates the effectiveness of our model in supporting the relation triple extraction.
At the bottom of Table 2, the NYT10 and NYT11 mainly contain relation instances belong to the Normal class. In comparison to robust model TPLinker which is a model based on a local semantic feature that utilizes four different tags for entity and relation detection, our model improves the performance about 0.5%, 0.6%, 1.3%, 1.4% in F1 scores, respectively. The improvement indicates that our model also has a better generalization capability. Such results confirm the effectiveness of employing the bi-consolidating module to extract all relational triples, as it enhances the refinement of semantic features by incorporating both local and global semantic features.
The improvements of the bi-consolidating model are mainly based on three advantages: first, it effectively enhances the local feature information from adjacent regions and mitigate noise in neighbouring neighbours. Secondly, it assimilates relevant information from a global perspective and learn remote semantic dependencies in a sentence, thereby effectively addressing the joint extraction task. Lastly, the utilization of two-dimensional sentence representation gives more semantic information to the sentences and effectively addresses the significant issue of semantic overlapping in relational triple extraction.
@@ 4.3 Ablation Study
The purpose of this experiment is to evaluate the contributions of each component in the bi-consolidating model. Each component is removed individually to evaluate its influence on the final performance. The dataset used for this experiment is the $\mathrm{WebNLG}$. The outcome is presented in Table 3.
Table 3: Ablation studies were performed on the WebNLG dataset using the bi-consolidating model under seven distinct settings.
@@
(3) w/o Local Consolidation denotes the model, which mainly uses global semantic features to enhance the semantic information. In this setting, the F1 score drops about 0.7% F1 score. The result serves as compelling evidence for the efficacy of the Global Consolidation model. This module enhances important features and attenuate unimportant ones controlling the scale, thus rendering the extracted features more directed. Besides, the outcome demonstrates that our model outperforms “Global Consolidation” by a marginal improvement of 0.2% in terms of the F1 score. It indicates its stronger performance in leveraging global features over local features.
(4) w/o Spatial Attention denotes to the model excluding the Spatial Attention Module from Global Consolidation, which results in the inadequate highlighting of the emphasis on “where”. The F1 score experiences a 0.3% decrease in comparison to the performance of the complete model, thus affirming the significance of the Spatial Attention Module. Spatial attention module allows the model to direct its attention towards salient keywords linked to entities or relationships within a given sentence, thereby reinforcing the significance of the respective word vectors.
(5) w/o Channel Attention denotes the model removing the Channel Attention Module from Global Consolidation. Under this condition, it fails to effectively highlight feature information about “what”. Our experiments showed that the model performed slightly lower than the “w/o Spatial Attention”, which indicates that channel attention module plays a vital role in Global Consolidation. The channel attention module acts as a core component by simulating the capture of global semantic features. It also indicates that the channel attention is applied globally, whereas the spatial attention operates on a local scale.
(6) w/o MaxPooling denotes the model without the MaxPooling operation in Global Consolidation. It hinders the model’s ability to focus on the most salient features within a given sentence. Our experimental results illustrate that the absence of MaxPooling operation has an adverse impact on both Spatial Attention Module and Channel Attention Module, ultimately leading to a decline of 0.4% in the F1 score of the model. This serves as proof that GAP can only acquire sub-optimal features due to the absence of crucial information. However, MaxPooling can effectively complement its shortcomings.
(7) w/o AvgPooling denotes the model without the AvgPooling operation in Global Consolidation. Our experiments indicate that this model performs slightly worse than "w/o MaxPooling", implying that average pooling removes some redundant information while maximizing retention of crucial information within a sentence. Therefore, AvgPooling represents the most efficient method for preserving the semantic features of a sentence.
@@ 4.4 Performance on Different Overlapping Patterns
Table 4: F1-score (%) on sentences with varying overlapping patterns and varying numbers of triples. $\mathrm{~{}I}$ represents the number of triples within a sentence. The superscript §indicates that Zheng et al. (2021) reported the results.
Dataset Model Normal $\mathrm{EPO}$ SEO SOO $\mathrm{I}=1$ $\mathrm{~{}I}=2$ $\mathrm{~{}I}=3$ $\mathrm{~{}I}=4$ $\mathrm{~{}I}\geq 5$
[…]
@@
[…]
@@
The result in Table 4 shows that the bi-consolidating model attains the highest F1-score across all these datasets. Particularly, the performance is improved considerably in SOO and $I\geq 5$ scenarios, where the overlapping problem is more serious.
The SOO scenario has two distinct characteristics. First, nested entities are difficult to identify, due to their shared contextual features within a sentence. For example, in the triple (“Percy Smith”, Family name, “Smith”), the object entity is a substring of the subject entity. Second, the difficulty arises when the subject and object contain the same words. For instance, the sentence “The ethnic group of native Americans are from the United States where the president is the leader and baked Alaska is a dish” in $\mathrm{WebNLG}^{*}$, which contains a triple (“States”, ethnicGroup, “States”). Furthermore, compared with related works, our model shows a significant improvement in $I\geq 5$, where a sentence may hold patterns of SEO, EPO, and SOO all at once. However, our model demonstrates superior performance on both datasets for SOO and $I\geq 5$. This serves as a clear evidence to show that our model effectively addresses the overlapping triplet problem and its robustness in handling these scenarios.
@@ 4.5 Performance on Different Sub-tasks
Previous studies (Sui et al., 2023) have shown that the joint extraction task faces two challenges: entity pair recognition and triple formation. For example, suppose a sentence containing three instances: (“Macron”, President of, “France”), (“Macron”, Live in, “France”), (“Macron”, Place of birth, “France”). The recall on relation $r$ may decrease to 0.67 if one relation is misclassified. On the other hand, the recall on entity pair $(s,o)$ may decline to 0 if one entity pair is mistakenly recognized.
@@
Comparing the performance between the $\mathrm{WebNLG}^{*}$ dataset and the $\mathrm{NYT}^{*}$ dataset, the number of EPO patterns in $\mathrm{WebNLG}^{*}$ is significantly lower than $\mathrm{NYT}^{*}$ (26 vs. 978). Furthermore, the dataset $\mathrm{WebNLG}^{*}$ is annotated with a greater number of relation types compared to $\mathrm{NYT}^{*}$ (171 vs. 24). Therefore, the relation triple extraction task on $\mathrm{WebNLG}^{*}$ is more challenging than on the $\mathrm{NYT}^{*}$ task. The result shows that the bi-consolidating model achieves robust performance on the two datasets. The results demonstrate the bi-consolidating model’s robust performance on both datasets.

### v2

@@ 3.3 Triple Generating Module
In this module, the 2D sentence representation $\mathbf{T}^{so}$ is fed into a two-layer perceptron to learn a table feature for each relation candidate. The table feature for a relation $r$ is denoted as $\mathbf{T}^{r}$. Each item in $\mathbf{T}^{r}$ represents the label feature for a token pair. Specifically, for a token pair $(x_{i},x_{j})$, we represent its label feature as $\mathbf{T}^{r}(i,j)$, which is computed by the following Equation (10).
@@
where $W\in\mathbb{R}^{D_{h}\times d}$, $b$ are trainable weight and bias. $W_{r}\in\mathbb{R}^{d\times K\times|Y|}$, $b_{r}$ are trainable weight and bias. Here, $K$ signifies the number of predefined relations, $|Y|$ denotes the number of predicted labels. The function $drop(\cdot)$ denotes the dropout strategy (Srivastava et al., 2014) utilized for mitigating overfitting.
Next, all relation triples can be generated by utilizing the table features $\mathbf{T}\in\mathbb{R}^{N\times N\times K\times|Y|}$. Specifically, for every relation, the method outlined in Equation (11) is employed to populate its corresponding table.
$\hat{\operatorname{table}_{r}}(i,j)=\operatorname{softmax}\left(\mathbf{T}^{r}(i,j)\right)$ (11)
@@
Finally, the filled tables are decoded. All triples are deduced using the following method. The subject can be obtained through splicing from “B-E” to “E-E”. The object can be obtained through splicing from “B-B” to “B-E”. They result in the acquisition of the relational triple. It is noteworthy that the two paired entities share a common “B-E”.
@@
$\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j=1}^{N}\log\left(\operatorname{table}_{r}(i,j)=\operatorname{gold}_{r}(i,j)\right)$
where $\operatorname{gold}_{r}(i,j)$ denotes the gold tag obtained from annotations. The codes to implement our bi-consolidating model is available at: https://github.com/xiaocen-luo/bi-consolidating-model.
@@ 4.1 Experimental Settings
Train Valid Test Relation Normal SEO EPO SOO $I=1$ $I=2$ $I=3$ $I=4$ $I\geq 5$ Triples
@@
In these datasets, both NYT and WebNLG have two distinct versions. Specifically, $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ are annotated with the tail tokens of a given entity. $\mathrm{NYT}$ and $\mathrm{WebNLG}$ are annotated with the whole range of an entity. Table 1 presents the statistical information of these datasets.
@@
To evaluate the performance of our models, we utilized Partial Match and Exact Match to measure the performance on the NYT, NYT10, NYT11 and WebNLG datasets. Partial Match considers a triple (subject, relation, object) to be correct only if the relation, as well as the tail tokens of both subject and object, is accurate. On the other hand, Exact Match deems a correct triple only if the entire span of both entities and the relation are precisely matched.
@@
Implementation Details: in order to ensure fair comparison, we utilize the Bert-base-cased model  [^1] Available at https://huggingface.co/bert-base-cased, consisting of 12 Transformer blocks with a hidden size of $D_{h}=768$. For the WebNLG dataset, a batch size of 6 is set, whereas for other datasets, the batch size is configured to be 5. We configure the local consolidation module as ‘$[C-A_{r}-R-V]$’ and optimize all parameters using the Adam (Kingma and Ba, 2014) with the learning rate of $1\times 10^{-5}$. In Equation (10), the hidden size of $d=3\times D_{h}$ is utilized for the NYT and $\mathrm{NYT}^{*}$ dataset, whereas for the remaining datasets, the hidden size of $d=2\times D_{h}$ is employed. Additionally, the maximum sentence length for input is set at 100.
@@ 4.2 Main Results
In this experiment, our model is compared with nineteen related works. In these models, GraphRel, NovalTagging, CopyRE, ETL-span, and MultiHead utilized LSTM networks as an encoder to generate token representations, while the others applied a pre-trained BERT model to do so. The performance is evaluated in terms of the Partial Match and the Exact Match. The result is presented in Table 2.
@@
Model Partial Match Exact Match
@@
PRGC (Zheng et al., 2021) $\underline{93.3}$ 91.9 92.6 94.0 92.1 93.0 $\mathbf{93.5}$ 91.9 92.7 89.9 87.2 88.5
@@
BiRTE (Ren et al., 2022) 92.2 $\underline{93.8}$ 93.0 93.2 94.0 93.6 91.9 $\mathbf{93.7}$ 92.8 89.0 89.5 89.3
@@
BTDM (Zhang et al., 2023b) 93.0 92.5 92.7 94.1 93.5 93.8 93.1 92.4 92.7 90.9 90.1 90.5
OneRel (Shang et al., 2022) 92.8 92.9 92.8 94.1 94.4 94.3 93.2 92.6 92.9 91.8 90.3 91.0
BitCoin (He et al., 2023) 92.9 92.8 93.1 $\underline{94.4}$ 94.5 94.4 93.1 92.6 $\underline{92.8}$ $\underline{91.9}$ $\underline{90.5}$ $\underline{91.2}$
UniRel (Tang et al., 2022) $\mathbf{93.5}$ $\mathbf{94.0}$ $\mathbf{93.7}$ $\mathbf{94.8}$ $\underline{94.6}$ $\underline{94.7}$
Ours 92.9 93.3 $\underline{93.1}$ 94.1 $\mathbf{96.0}$ $\mathbf{95.0}$ 93.1 $\underline{92.8}$ $\mathbf{92.9}$ $\mathbf{92.2}$ $\mathbf{91.3}$ $\mathbf{91.7}$
Model Partial Match Exact Match
@@
As Table 2 shown, compared with other models, our model exhibits competitive performance. In the $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ datasets, relation triples are solely annotated with the final word of a given entity. The improvement of our model suggests that the bi-consolidating model is successful in learning global features relevant to a relation instance. On the other hand, the improvement in the $\mathrm{NYT}$ and $\mathrm{WebNLG}$ also indicates that the bi-consolidating model has the ability to correctly extract the whole span of entities, which heavily depends local features of a sentence. In the $\mathrm{NYT}^{*}$ dataset with Partial Match, the UniRel model has the best performance. The reason may be that the UniRel model jointly encodes the representations of entities and relations within a concatenated natural language sequence, which is effective to improve the Partial Match measurement. Except the UniRel model, our model achieves the best performance in another three datasets. The outcome demonstrates the effectiveness of our model in supporting the relation triple extraction.
At the bottom of Table 2, the NYT10 and NYT11 mainly contain relation instances belong to the Normal class. In comparison to robust model TPLinker which is a model based on a local semantic feature that utilizes four different tags for entity and relation detection, our model improves the performance about 0.5%, 0.6%, 1.3%, 1.4% in F1 scores, respectively. The improvement indicates that our model also has a better generalization capability. Such results confirm the effectiveness of employing the bi-consolidating module to extract all relational triples. It enhances the refinement of semantic features by incorporating both local and global semantic features.
The improvements of the bi-consolidating model are mainly based on three advantages. First, it effectively enhances the local feature information from adjacent regions and mitigate noise in neighbouring neighbours. Secondly, it assimilates relevant information from a global perspective and learn remote semantic dependencies in a sentence, thereby effectively addressing the joint extraction task. Lastly, the utilization of two-dimensional sentence representation gives more semantic information to the sentences and effectively addresses the significant issue of semantic overlapping in relational triple extraction.
@@ 4.3 Ablation Study
The purpose of this experiment is to evaluate the contributions of each component in the bi-consolidating model. Each component is removed individually to evaluate its influence on the final performance. The position embedding and attention embedding in Equation (3) are also independently removed to show the influence of sentence encoding. The dataset used for this experiment is the $\mathrm{WebNLG}$. The outcome is presented in Table 3. The influence of each component is discussed as follows.
Table 3: Ablation studies were performed on the WebNLG dataset using the bi-consolidating model under ten distinct settings.
@@
(3) w/o Local Consolidation denotes the model, which mainly uses global semantic features to enhance the semantic information. In this setting, the F1 score drops about 0.7% F1 score. The result serves as compelling evidence for the efficacy of the Global Consolidation model. This module enhances important features and attenuate unimportant ones controlling the scale, thus rendering the extracted features more directed. Besides, the outcome demonstrates that our model outperforms “w/o Global Consolidation” by a marginal improvement of 0.2% in terms of the F1 score. It indicates its stronger performance in leveraging global features over local features.
(4) w/o Spatial Attention denotes the model excluding the Spatial Attention Module from Global Consolidation, which results in a less prominent emphasis on salient keywords associated with entities or relations in a given sentence. The F1 score experiences a 0.3% decrease in comparison to the performance of the complete model, thus affirming the significance of the Spatial Attention Module. Spatial attention module allows the model to direct its attention towards salient keywords linked to entities or relationships within a given sentence, thereby reinforcing the significance of the respective word vectors.
(5) w/o Channel Attention denotes the model removing the Channel Attention Module from Global Consolidation. Under this condition, it fails to effectively highlight global semantic features. Our experiments showed that the model performed slightly lower than the “w/o Spatial Attention”, which indicates that channel attention module plays a vital role in Global Consolidation. The channel attention module acts as a core component by simulating the capture of global semantic features. It also indicates that the channel attention is applied globally, whereas the spatial attention operates on a local scale.
(6) w/o MaxPooling denotes the model without the MaxPooling operation in Global Consolidation. It hinders the model’s ability to focus on the most salient features within a given sentence. Our experimental results illustrate that the absence of the MaxPooling operation has an adverse impact on both Spatial Attention Module and Channel Attention Module, ultimately leading to a decline of 0.4% in the F1 score of the model. This serves as proof that GAP can only acquire sub-optimal features due to the absence of crucial information. However, MaxPooling can effectively complement its shortcomings.
(7) w/o AvgPooling denotes the model without the AvgPooling operation in Global Consolidation. Our experiments indicate that this model performs slightly worse than “w/o MaxPooling”, implying that average pooling removes some redundant information while maximizing retention of crucial information within a sentence. Therefore, AvgPooling represents the most efficient method for preserving the semantic features of a sentence.
@@
(9) w/o Position Embedding: the purpose of position embedding is to capture the relative positional information between each pair of tokens within a sentence. When the Position Embedding is omitted from the sentence encoding model, the performance decreases, which suggest that the absence of relative location information for each pair of tokens is influential on the final performance.
(10) w/o Attention Embedding: the attention embedding is derived from the raw input using an attention layer. This embedding is learned by assigning weights to input features based on their relevance or importance to the output. Experimental results indicate that the presence of Attention Embedding positively influences the learning process by refining two-dimensional representations.
@@ 4.4 Performance on Different Overlapping Patterns
Table 4: Performance on sentences with varying overlapping patterns and varying numbers of triples. The superscript §  indicates that the result is reported by Zheng et al. (2021).
Dataset Model Normal $\mathrm{EPO}$ $\mathrm{SEO}$ $\mathrm{SOO}$ $\mathrm{I}=1$ $\mathrm{~{}I}=2$ $\mathrm{~{}I}=3$ $\mathrm{~{}I}=4$ $\mathrm{~{}I}\geq 5$
$\mathrm{NYT}^{*}$ CasRel 87.3 92.0 91.4 $\text{77.0}^{\S}$ 88.2 90.3 91.9 94.2 83.7
TPLinker 90.1 94.0 93.4 $\text{90.1}^{\S}$ 90.0 92.8 93.1 96.1 90.0
@@
$\mathrm{WebNLG}^{*}$ CasRel 89.4 94.7 92.2 $\text{90.4}^{\S}$ 89.3 90.8 94.2 92.4 90.9
TPLinker 87.9 95.3 92.5 $\text{86.0}^{\S}$ 88.0 90.1 94.6 93.3 91.6
@@
The result in Table 4 shows that the bi-consolidating model achieves the highest F1-score across all these datasets. Particularly, the performance is improved considerably in $\mathrm{SOO}$ and $I\geq 5$ scenarios, where the overlapping problem is more serious.
The SOO scenario has two distinct characteristics. First, nested entities are difficult to identify, due to their shared contextual features within a sentence. For example, in the triple (“Percy Smith”, “Family name”, “Smith”), the object entity is a substring of the subject entity. Second, the difficulty arises when the subject and object contain the same words. For instance, the sentence “The ethnic group of native Americans are from the United States where the president is the leader and baked Alaska is a dish” in $\mathrm{WebNLG}^{*}$, which contains a triple (“States”, “ethnicGroup”, “States”). Furthermore, compared with related works, our model shows a significant improvement in $I\geq 5$, where a sentence may hold patterns of SEO, EPO, and SOO all at once. This serves as a clear evidence to show that our model effectively addresses the overlapping triplet problem and its robustness in handling these scenarios.
@@ 4.5 Performance on Different Sub-tasks
Previous studies (Sui et al., 2023) have shown that the joint extraction task faces two challenges: entity pair recognition and triple formation. For example, suppose a sentence containing three instances: (“Macron”, “President of”, “France”), (“Macron”, “Live in”, “France”), (“Macron”, “Place of birth”, “France”). The recall on relation $r$ may decrease to 0.67 if one relation is misclassified. On the other hand, the recall on entity pair $(s,o)$ may decline to 0 if one entity pair is mistakenly recognized.
@@
When comparing the features of the $\mathrm{WebNLG}^{*}$ dataset and the $\mathrm{NYT}^{*}$ dataset, it is observed that the number of training sets in $\mathrm{WebNLG}^{*}$ is significantly lower than that in $\mathrm{NYT}^{*}$ (5019 vs. 56195). Additionally, $\mathrm{WebNLG}^{*}$ contains a higher number of predefined relations (171 vs. 24) and nearly twice as many SOO patterns as $\mathrm{NYT}^{*}$ (84 vs. 45). Therefore, the relation triple extraction task on $\mathrm{WebNLG}^{*}$ is more challenging than on the $\mathrm{NYT}^{*}$ task. The result shows that the bi-consolidating model achieves robust performance on the two datasets. The results demonstrate the bi-consolidating model’s robust performance on both datasets.